OpenStudy (anonymous):

Use Lagrange multipliers to find the shortest distance between the line (2,4,4)+s(4,1,5) and (1,-3,2)+t(-2,3,1).

OpenStudy (anonymous):

specifically, I don't know how to work with the equations of the lines in that form.

OpenStudy (anonymous):

I know I will use the distance squared, with the constraints as those two lines, but the equation of the lines is not in a form I can work with...

OpenStudy (anonymous):

What is a form that you can work with? (2,4,4)+s(4,1,5) means the line joining the points (2,4,4) and (6,5,9). ( (6,5,9) = (2,4,4)+ (4,1,5) ).

OpenStudy (anonymous):

What I mean is that the lines are not in the form like ax+by=cz, I don't know how to work with the parametric equations of the lines

OpenStudy (helder_edwin):

u have the following sets \[ \large L_1=\{(2,4,4)+s(4,1,5):s\in\mathbb{R}\} \] and \[ \large L_2=\{(1,-3,2)+t(-2,3,1):t\in\mathbb{R}\} \] then the distance between any couple of points of these sets would be \[ \large\sqrt{(4s+2t+1)^2+(s-3t+7)^2+(5s-t+2)^2}=d(s,t) \] this is the function u have to minimize.

OpenStudy (anonymous):

thanks :D

OpenStudy (helder_edwin):

u r welcome